### v2

# How Useful is Continued Pre-Training for Generative Unsupervised Domain Adaptation?

Abstract
Recent breakthroughs in scale have enabled the emergence of powerful generative language models, and the ability to fine-tune these models on various tasks by casting them into prompts or instructions. In this landscape, the problem of Unsupervised Domain Adaptation (UDA), or the problem of leveraging knowledge from a labeled source domain to an unlabeled target domain, has been left behind, with recent UDA methods still addressing discriminative classification. In particular, two popular UDA approaches, involving Continued Pre-Training (CPT) and learning domain invariant representations, have been under-explored in the generative setting, signaling a gap. In this work, we evaluate the utility of CPT for generative UDA. We first perform an empirical evaluation to measure the trade-offs between CPT and strong methods promoting domain invariance. We further evaluate how well the benefits of CPT extend to different architectures, tuning methods and data regimes. We then motivate the use of CPT by studying to what degree it benefits classification performance on the target domain. Finally, we attempt to understand the mechanism behind which CPT improves classification performance on the unlabeled target domain. Our findings suggest that a implicitly learns the downstream task while predicting masked words informative to that task. Our work connects the body of UDA research with that of instruction tuning, enabling an initial step towards a wider applicability of modern language models.
1 Introduction
Recent advancements in the pre-training of language models have enabled the widespread use of powerful generative models, which can be leveraged across multiple domains with no training (Brown et al., 2020; Scao et al., 2022; Touvron et al., 2023, inter alia). Despite these advancements, these autoregressive models are still fragile under certain kinds of data distribution shifts, making their applications across these domains challenging (Ribeiro et al., 2020). This is addressed, in part, by the concept of instruction tuning with templates (Zhang et al., 2023; Sanh et al., 2022; Ouyang et al., 2022; Wang et al., 2022; Wei et al., 2022), enabling the learning of new tasks without any randomly initialized parameters.
The problem of unsupervised domain adaptation (UDA) leverages learned knowledge from a labeled source domain to an unlabeled target domain (Pan & Yang, 2010; Ganin & Lempitsky, 2015; Long et al., 2015, inter alia). It is useful for adaptation to unlabeled domains with high labeling costs, where supervised instruction tuning does not suffice. Despite the pervasive need for models to generalize to such domains, recent UDA methods still address discriminative classification, barring the application of these approaches to recent generative models. In particular, Continued Pre-Training and Domain Invariance based methods, two widely popular classes of UDA approaches (Ramponi & Plank, 2020), are completely unexplored in the generative setting.
The Continued Pre-Training (CPT) approach involves extended pre-training on a domain or task, followed by supervised training on the downstream task (Gururangan et al., 2020). This approach has been widely used for adaptation to labeled domains (Gao et al., 2021; Kim et al., 2021; Hung et al., 2023, inter alia), and in the UDA setup for unlabeled domains (Han & Eisenstein, 2019; Zhang et al., 2021b; Karouzos et al., 2021; Pfeiffer et al., 2020; Parović et al., 2023). Invariance based approaches attempt to learn representations that are invariant across domains (Tzeng et al., 2014; Ganin et al., 2016; Wu & Shi, 2022; Guo et al., 2022), with the notion that when the learned representations from both domains cannot be distinguished by a classifier and the classifier performs well on the source domain, it will also exhibit strong performance on the target domain. These two classes of methods introduce a trade-off: invariance based methods suffer from instability issues (Han & Eisenstein, 2019; Kashyap et al., 2021), while continued pre-training requires a larger computational budget. But how would this trade-off extrapolate to the generative setting? For example, invariance based methods are well motivated in discriminative tasks, where there is a clear decision boundary; however, the same does not hold for generative tasks.
To address these gaps, we introduce the setting of Generative UDA, where an autoregressive model is trained to leverage knowledge from a labeled source domain to an unlabeled target domain, using only next word prediction as its objective. We formalize the use of CPT for this setting in Section 2, and then attempt to explore and understand the behaviour of CPT for Generative UDA. We begin by performing an empirical analysis on 40 real-world domain pairs to explore the tradeoff between continued pre-training and invariance based approaches, and find vanilla CPT to be competitive with and significantly more stable than a state of the art invariance based approach (Section 3). We then stress test CPT, by applying it across varying model architectures and scales, tuning approaches and data regimes. We find that CPT is robust across these settings, unlike our invariance based approach (Section 4).
With recent language models being trained across vast corpora which may include domains similar to the target domain, the requirement for continued pre-training may be raised to question. In Section 5, we show that continued pre-training is indeed essential for strong downstream performance on the target domain, and this performance rapidly degrades with limited target domain exposure. Finally, we attempt to shed light on how masking plays a role in improving classification accuracy on the unlabeled target domain in Section 6. We find that the model may implicitly learn the downstream task as it predicts masked words that are informative to the downstream task.
Our work attempts to connect the body of UDA research with recent trends in language modeling, by providing a set of insights into the behaviour of the popular class of continued pre-training approaches, in the Generative UDA setting. We hope this enables an initial step towards a wider applicability of modern language models.
2 Continued Pre-Training for Generative UDA
2.1 Preliminaries: The UDA Problem
We consider a text classification task, where $\mathcal{X}$ is the input space of all text sentences and $\mathcal{Y}=\{1,...K\}$ is the label space. In the UDA problem, we have access to a source labeled dataset $\mathcal{D}_{\text{src}}=\{(x_{i},y_{i})\}_{i=1}^{N}$ consisting of samples from a joint distribution $P_{\text{src}}$, and a target unlabeled dataset $\mathcal{D}_{\text{tgt}}=\{x_{j}\}_{j=1}^{M}$ sampling from a target input distribution $P^{\mathcal{X}}_{\text{tgt}}$. We further denote $P^{\mathcal{X}}_{\text{src}}$ as the marginal distribution of $P_{\text{src}}$ on $\mathcal{X}$, where $P^{\mathcal{X}}_{\text{src}}\neq P^{\mathcal{X}}_{\text{tgt}}$. The goal of UDA is to learn a function $f:\mathcal{X}\to\mathcal{Y}$ that minimizes the error rate $\mathbb{E}_{x\sim P^{\mathcal{X}}_{\text{tgt}}}\;\mathbbm{1}[f(x)\neq y]$.
2.2 CPT for UDA as a Sequence of Prompt Based Tasks
We now formalize the extension of continued pre-training to the setting of generative UDA. We use the traditional two-phase training pipeline from Gururangan et al. (2020) [^1] While we use the two-phase multi-task training pipeline (sequential) in our main experiments, in Appendix I, we show that an equivalent single-phase multi-task training pipeline (joint) results in similar performance.. The first phase uses templates to cast the source and target domain sequences into an autoregressive pre-training task. [^2] We investigate mask language modeling for T5 models and switch to causal language modeling for decoder-only models with a few simple template changes. We compare both in Section 4. The second phase applies supervised instruction-tuning of the model on source-labeled data.
Task 1: Autoregressive Continued Pre-training
We reuse the input sequences from the source-labeled dataset $\mathcal{D}_{\text{src}}$ as the source-unlabeled dataset, denoted as $\mathcal{D}_{\text{src}}^{\mathcal{X}}$. Next, similar to Raffel et al. (2020); Song et al. (2019), for an unlabeled sequence $x\in\mathcal{D}_{\text{src}}^{\mathcal{X}}$ and $\mathcal{D}_{\text{src}}$, we use a prompt template to convert the sequence $x$ to an input-output sequence pair, i.e., $\mathbb{M}(x)=(\tilde{x},\tilde{y})$. For masked language modeling (MLM), an instruction is pre-pended to a randomly masked sequence $x$ to create $\tilde{x}$. The output sequence $\tilde{y}$ is a concatenation of masked words from $x$. For example, given $x=$“The movie was so cool! Two hours of fun.”, we construct $\tilde{x}=$“Fill in the blanks: ”The _ cool! Two hours _”, and $\tilde{y}=$ <sep> movie was so <sep> of fun. <sep>. For causal language modeling (CLM), $\tilde{x}=x$.
Given $(\tilde{x},\tilde{y})$, we train an autoregressive LM parameterized by $\theta$ to minimize the negative log-likelihood loss averaged over output words and the total loss over a corpus $\mathcal{D}=\mathcal{D}_{\text{src}}^{\mathcal{X}}\cup\mathcal{D}_{\text{tgt}}$.
$\ell(\tilde{x},\tilde{y};\theta)=-\frac{1}{|\tilde{y}|}\sum_{t}\log P_{\theta}(\tilde{y}_{t}|\tilde{x},\tilde{y}_{1:t-1});~{}~{}~{}~{}~{}~{}\mathcal{L}_{\text{CPT}}(\mathcal{D};\theta)={\frac{1}{|\mathcal{D}|}}\sum_{x\in\mathcal{D}}\ell(\mathbb{M}(x);\theta).$ (1)
Task 2: Source Supervised Instruction-tuning
In the second phase, we use labeled data from the source domain to train the model on the downstream classification task. Similar to the first phase, we use prompts [^3] Prompt templates were selected from the Public Pool of Prompts (Bach et al., 2022). to generate input-output sequence pairs: $\mathbb{C}(x,y)=(\tilde{x},\tilde{y})\;\forall\;(x,y)\in\mathcal{D}_{\text{src}}$. For example, for sentiment classification, if $x=$“I like this movie.”, $y=1\Rightarrow$ $\tilde{x}=$ “[$x$] Is this sentence positive or negative?”, $\tilde{y}=$ “Positive”.
Given the augmented sequence pair $(\tilde{x},\tilde{y})$ and the model trained after the first phase, we compute the same negative log-likelihood loss $\ell(\tilde{x},\tilde{y};\theta)$ in Eq. (1). Finally, we define the total loss on the source-labeled dataset in the second phase as:
$\mathcal{L}_{\text{CLS}}(\mathcal{D}_{\text{src}};\theta)={\frac{1}{|\mathcal{D}_{\text{src}}|}}\sum_{(x,y)\in\mathcal{D}_{\text{src}}}l(\mathbb{C}(x,y);\theta).$ (2)
After training, we follow the practice of Liu et al. (2022) to convert a label string $\tilde{y}$ to its corresponding label $y$ at test time for evaluation.
3 Evaluating the Efficacy of Continued Pre-training for Generative UDA
3.1 Experimental Setup
Datasets
We use the MNLI and Amazon Review classification datasets, which are widely used UDA benchmarks (Malik et al., 2023; Karouzos et al., 2021; Guo et al., 2020). The MNLI corpus (Williams et al., 2018) contains sentence pairs across five genres: Travel (T), Fiction (F), Government (G), Slate (S), and Telephone (Te). The task classifies every sentence pair as entailment, neutral, or contradiction. The Multi-Domain Sentiment Analysis Dataset (Blitzer et al., 2007) contains binary sentiment reviews for different types of Amazon products. We use reviews from the Apparel (A), Baby (B), Books (Bo), Cameras (C), and Movies (M) domains. We evaluate a total of 40 pairs of source and target domains, across the two datasets. Appendix A contains more details about the datasets.
Models and Tuning Methods
Our main experiments use the T5v1.1 base model and (IA)3 (Liu et al., 2022) PEFT method. T5v1.1 is an improved version of the original T5 model (Raffel et al., 2020), and unlike the original T5 model, it is not trained on any supervised datasets. We then extend our evaluation to different model architectures (T0, GPT-2), tuning methods (full fine-tuning, adapters) and data regimes (Section 4).
Training
Each training phase is 30,000 steps long for MNLI and 15,000 steps for the Amazon dataset. We use Adam, a batch size of 8, and a learning rate of 0.003. We set the maximum sequence length to 256 tokens. We use length normalization during evaluation, as proposed by Liu et al. (2022). For each experiment, we report the mean and standard deviation across 3 runs. More details can be found in Appendix B.
Baselines
Since our goal is to study the behaviour of CPT for generative UDA, we it compare with a simple supervised baseline, and a state-of-the-art invariance based approach.
  • Src+Tgt (All labeled): We fine-tune the model using labeled data from both the source and target domains. This serves as an upper bound on target domain performance.
  • UDAPTER: Malik et al. (2023) propose an invariance based method that measures the multi-kernel maximum mean discrepancy (MK-MMD) (Gretton et al., 2012; Bousmalis et al., 2016) between source and target embeddings from each transformer layer and sums them to obtain an aggregate loss $\mathcal{L}_{\text{div}}$. The final loss is the weighted sum of $\mathcal{L}_{\text{div}}$ and the classification loss, i.e., $\mathcal{L}=\lambda\;\mathcal{L}_{\text{cls}}+(1-\lambda)\;\mathcal{L}_{\text{div}}$, where $\lambda$ gradually changes from 0 to 1 during training. Here, we use the embeddings from a model as it is being instruction tuned on the downstream classification task. Their method achieves state-of-the-art performance, and outperforms popular UDA approaches like DANN (Ganin et al., 2016) and DSN (Bousmalis et al., 2016); thus we only compare CPT with this approach.
3.2 Continued Pre-training is Competitive with Domain-Invariance Methods
Dataset Src+Tgt UDAPTER CPT
Amazon 92.66 (0.45) 89.02 (2.17) 89.34 (0.48)
MNLI 78.14 (0.25) 70.19 (1.71) 74.12 (0.68)
Table 1: Avg. target-domain classification accuracy and standard deviation over 3 runs on Amazon Review and MNLI.
Performance
We compare CPT with other baselines over 40 domain pairs of the MNLI and Amazon Review datasets, and report the average accuracies over all pairs in Table 1. We see that CPT is competitive to UDAPTER. (Appendix C contains detailed results over 40 pairs and significance tests to check for competitiveness.) Interestingly, a visualization of sentence embeddings in Figure 8 (Appendix C) suggests that representations learned through CPT are not domain invariant. In addition to the MMD based method of Malik et al. (2023), we also compare CPT on one domain pair with other methods that promote domain invariance (DANN (Ganin et al., 2016), CORAL (Sun et al., 2017)) and weight interpolation (Ilharco et al., 2022) in Appendix C, further confirming the competitiveness of CPT.
Stability
CPT performs more stably than UDAPTER, with the invariance-based method often reporting a variance of over 20% across runs (Table 6 in Appendix C). For example, for the MNLI pair Fiction (F) $\to$ Government (G), minimizing UDAPTER yields a variance of 23.4% across runs. This observation is consistent with existing findings (Kashyap et al., 2021; Han & Eisenstein, 2019) that minimizing divergence measures like MMD, when combined with auxiliary task-specific loss functions, result in training instabilities and vanishing gradients. We discuss this in more detail in Appendix J.
4 How General are the Benefits of CPT?
[…]
CPT Helps Decoder-only Models.
We extend our analysis from MLM with encoder-decoder language models, to causal language modeling (CLM) with decoder-only language models, using GPT-2 (medium) (Radford et al., 2019). We perform CLM in the first training phase by simply training the model for next-word prediction given the original sequence. Table 2 shows that CPT provides strong improvements on the target domain in comparison to the invariance based baseline.
CPT Outweighs Invariance-based Methods for Instruction-tuned Models.
We evaluate the performance of CPT over T5v1.1 XL (3B parameters) and the instruction-tuned T0 (3B parameters) (Sanh et al., 2022). Figure 2 (Table 8 of Appendix D) shows a wider gap between UDAPTER and CPT with higher model capacity, and this gap is further increased with instruction tuning. We hypothesize that this gap is due to the vast difference between the objectives of domain invariance and instruction tuning.
Method Accuracy
Src+Tgt 79.8 (0.3)
UDAPTER 66.0 (1.1)
CPT 75.8 (0.4)
Table 2: Performance of CPT with causal language modeling for the decoder-only GPT-2 model. CPT significantly outperforms the invariance-based method.
Figure 1: The performance gap between CPT and UDAPTER increases with larger models, from T5v1.1 Base (60M parameters) to T5v1.1 XL (3B parameters), and further increases with instruction tuning (T0 3B).
Figure 2: Performance of CPT across different tuning approaches with the T5v1.1 base model. CPT remains more powerful than UDAPTER across all tuning approaches.
CPT Benefits are Consistent across Tuning Approaches.
PEFT approaches have been shown to introduce resilience to domain shift (Fu et al., 2023). To isolate this effect from the CPT framework, we use T5v1.1 to evaluate CPT in a full fine-tuning setup. Additionally, we compare CPT with two PEFT approaches [^4] We choose Adapters because He et al. (2022) present a unified view of PEFT approaches which shows that the operations applied by Adapters are very similar to those of Prefix Tuning (Li & Liang, 2021) and LoRA (Hu et al., 2022). We choose (IA)3 since it is a state-of-the-art PEFT approach that uses a fraction of the learnable parameters of Adapters (More in Appendix E). : Adapters (Houlsby et al., 2019) and (IA)3 (Liu et al., 2022). We see in Figure 2 (Table 9 in Appendix E) that CPT continues to perform stronger than the domain invariance based UDAPTER method.
CPT Outperforms Invariance-based Methods in Low-data Regimes.
In this low-data experiment, we assume access to $k$ labeled source-domain examples. For CPT, we assume access to the full unlabeled dataset in both domains for the first training phase, and $k$-shot access to labeled source-domain examples for the second phase of supervised training. For a fair comparison, we also introduce a two-phase version of the UDAPTER pipeline—the first phase minimizes MMD between unlabeled source and target domain embeddings (full data access), while the second phase optimizes supervised training on the source domain ($k$-shot). Figure 4 (Table 10 in Appendix F) showcases CPT clearly outperforming both variants of UDAPTER, across three different models for $k=256$. Furthermore, Figure 4 (Table 11 in Appendix F) shows CPT providing consistent improvements in as low as 32-shots, unlike the unstable invariance-based approach.
Figure 3: Few-shot performance of CPT, for varying $k$. Unlike both variants of UDAPTER, CPT is stable and provides consistent benefits across number of shots.
Figure 4: Performance of CPT across different models, in a 256-shot learning setup. Unlike both variants of UDAPTER, CPT is stable and provides consistent benefits across models.
5 To What Extent is Target-Domain Exposure Beneficial?
Given the vast distributions language models are pre-trained on, a natural assumption might be that the model has already been exposed to a domain similar to the target domain during pre-training. This would mean that the model could simply extrapolate the learned downstream task from the source to the target domain, questioning the need for CPT.
[…]
Figure 5: Impact of Masking Rate on CPT. With high masking rates, the performance on the source domain is largely maintained, but the performance on the target domain rapidly deteriorates.
For a more fine-grained analysis, we investigate the impact of degree of exposure to the target domain, by varying the masking rates during the first phase of training. While masking 15% of a sequence is considered standard for random masking, previous work has shown that BERT-sized models (Devlin et al., 2019) can learn from as high as 80% masking rates during pre-training followed by adaptation to a labeled task (Wettig et al., 2023). The source-domain performance shown in Figure 5 (Table 12 in Appendix) matches this trend. However, high masking rates effectively reduce the exposure of the model to target data, strongly deteriorating the performance on the target domain [^5] With masking rates under the optimal value of 15%, the semantic and background features learned through model prediction of masked words is limited, hurting performance on the target domain.. We hypothesize that since the model never sees any labeled data of the target domain, it heavily depends on the signal it gets from the unlabeled data through masking.
Figure 6: Vocabulary overlap between label-informative words of the source and target domains. The numbers in the Venn diagram indicate the number of words in both sets.
Phase 1 Data Accuracy
Source Target
Source Only 93.3 (0.1) 76.5 (0.2)
Target Only 92.9 (0.4) 82.3 (0.7)
Source + Target 93.5 (0.4) 83.3 (0.5)
Table 3: Comparison of CPT with varying data exposure during the first phase of training. Performance on the target domain strongly benefits more from exposure to target domain, and is boosted further with exposure to the source domain.
6 Why does Word Prediction Aid Classification for Generative UDA?
In this section, we examine why predicting masked words of the source and target domains through CPT boosts sentence classification on the unlabeled target domain for generative UDA. We hypothesize that by having to predict masked words that are informative to the downstream task during pre-training, the model implicitly learns information about the downstream task. For example, given the masked sentence, “I really _ the movie, it was a fascinating watch.”, the masked word is indicative of the downstream task, in this case sentiment analysis. The model can only predict this masked word (which would be a positive word like “loved” or “enjoyed”) by using other words informative to the task (“fascinating”). Through this process, the model is essentially learning features which are useful to the downstream task, despite having no direct supervision.
To test this hypothesis, we quantize the “informativeness” of each word to a classification task: an informative word is highly correlated with any of the labels in the downstream task. [^6] These informative words are similar to pivot features (Blitzer et al., 2006; Ziser & Reichart, 2018; Ben-David et al., 2020, inter alia), with the exception that they are chosen based on information from the source domain only. Specifically, we follow Gururangan et al. (2018) and use pointwise mutual information (PMI) (Fano, 1961) of the word with respect to the class label:
$\text{PMI}(\text{word},\text{class})=\log\frac{p(\text{word},\text{class})}{p(\text{word})p(\text{class})},$
where we count the frequency of a word-class pair on $\mathcal{D}_{\text{src}}$ to estimate $p(\text{word},\text{class})$, and similarly count a word and a class individually on $\mathcal{D}_{\text{src}}$ to estimate $p(\text{word})$ and $p(\text{class})$.
[…]
How Masking helps learn Classification
We first confirm that label-informative words indeed impact the classification performance of the CPT model. We do this by masking informative words from a sentence at inference. Figure 7 (a) shows us that the performance of the model on the source domain is not impacted by masking uninformative words, but drops on masking informative words. However, how do we know how much of this bias towards label informative words was learned during the continued pre-training phase, rather than during supervised fine-tuning? To attempt to disentangle the impact of the training phases, we train the model through selective masking (informative or uninformative) in the first phase of training, and minimize the impact of the second phase by making it a few-shot task. Figure 7 (c) shows us that the model performs best on classification when trained to predict label informative words during masking, indicating that the model does indeed learn features relevant to the downstream task during the first phase of training.
The interplay between CPT and Classification for Generative UDA
We now extend this analysis to the target domain to understand how CPT plays a role in learning features from the unlabeled domain. Figure 7 (d) shows us that informative masking outperforms uninformative masking by a significant gap, once again signaling that the masking process helps the model implicitly learn the downstream task. However, unlike with the source domain, random masking results in the strongest performance. This is due to the domain mismatch: the informative words for the source and target domains are not identical (Figure 6), and the supervised training on the source domain adds a bias towards source-informative words. The mixture of these two sets of words are best predicted through random masking, explaining its strong performance.
This phenomenon also draws the observation that random masking is preferred to selective masking for generative UDA, contrary to single domain settings where informative masking is more useful (Levine et al., 2021; Gu et al., 2020).
Refer to caption Refer to caption
Figure 7: The impact of selective masking on classification performance of a CPT trained model. Left: Masking label informative words during inference degrades classification performance, compared to the original unmasked sequence. However, the removal of uninformative words does not impact the model on the source domain. Right: The impact of continued pre-training can be partially disentangled from that of supervised training by making the supervised training phase few-shot and training the model to mask informative or uninformative words during the continued pre-training phase. Informative masking is most beneficial for the source domain, indicating that the model learns task relevant features during masking. On the target domain, informative masking still captures some knowledge about the downstream task, however, the supervised training phase adds a bias towards source label informative words. Thus, random masking is most powerful.
7 Discussion
The Computational Trade-off of CPT
Our results in Section 3.2 show that continued pre-training and methods promoting domain invariance are competitive with each other. Continued pre-training suffers from the computational drawback of requiring an additional phase of training. Conversely, invariance based methods are difficult to optimize, possibly requiring more runs to achieve a stable optimum, and having a higher amortized computational cost. Inspired by Karouzos et al. (2021), we introduce a simple single phase variant of continued pre-training which is equivalent in performance to its two phase variant, nullifying the additional computational overhead of the approach (more details in Appendix I).
UDA in the age of LLMs
Recent breakthroughs in scale have showcased that large language models (LLMs) are highly powerful, and can perform various downstream tasks with limited or no training. This may raise question on the relevance of the UDA problem as a whole — does a model even require expensive adaptation to a domain it may have already been exposed to during pre-training? In addition to our analysis in Section 5, we argue that the requirement to adapt small models to unseen domains still holds in specific cases. Small supervised models have been shown to outperform zero-shot general-purpose LLMs on various downstream tasks (Huang et al., 2023; Zhu et al., 2023), and safety critical domains like healthcare and finance would benefit more from these models than a generalist LLM. Our study does not address how to better adapt to domains, rather we investigate ways a model may adapt to data unseen during pre-training. This is a question that holds for current LLMs, and will continue to hold as long as models are unable to access infinite data during pre-training.
8 Related Work
UDA through Promoting Domain Invariance
A major class of approaches in Model-centric UDA methods (Ramponi & Plank, 2020) aims to minimize $\mathcal{H}\Delta\mathcal{H}$ divergence (Ben-David et al., 2010) between the source and target domain features, through adversarial training (Tzeng et al., 2014; Ganin et al., 2016; Tzeng et al., 2017; Guo et al., 2022, inter alia) or through minimizing measures of domain similarity (Bousmalis et al., 2016; Ge et al., 2023). Malik et al. (2023) have shown the minimization of MMD to outperform other invariance based methods. However, past work has shown that domain-invariance is a weak constraint for adaptation (Zhao et al., 2019; Karouzos et al., 2021), could introduce domain-specific hyperparameters (Trung et al., 2022), and is also prone to instability issues (Han & Eisenstein, 2019; Sun et al., 2019; Wilson & Cook, 2020; Kashyap et al., 2021).
UDA through Continued Pre-Training
The limitations of invariance-based model-centric methods have encouraged the emergence of alternate approaches, based on self-supervised learning through contrastive learning (Kumar et al., 2022; Shen et al., 2022; Long et al., 2022), pseudo-labeling (Zhou & Li, 2005; Ruder & Plank, 2017, inter alia) or language model pre-training. Despite not being directly useful to certain downstream tasks (Uppaal et al., 2023), CPT has been used for adaptation to labeled tasks, in both full fine-tuning (Gururangan et al., 2020; Lee et al., 2020; Gao et al., 2021, inter alia) and PEFT setups (Kim et al., 2021; Hung et al., 2023). A smaller body of work has explored the utility of CPT in a UDA setup (Han & Eisenstein, 2019; Zhang et al., 2021b; Karouzos et al., 2021; Parović et al., 2023), identifying the class of methods to be more stable than invariance-based methods.
Generative UDA
The emergence of large language models (Brown et al., 2020; Scao et al., 2022; Touvron et al., 2023, inter alia) introduced the concept of instruction tuning with templates (Zhang et al., 2023; Sanh et al., 2022; Ouyang et al., 2022; Wang et al., 2022; Wei et al., 2022; Gao et al., 2021; Liu et al., 2023), enabling multi-task training without any task specific architectural changes. However, the framework of casting discriminative classification tasks into generative next word prediction tasks has not yet been extended to UDA. The closest work to this setting (Ben-David et al., 2021) uses a generative model to create domain identifier prompts and feed them back into the model, however the final task label prediction is still discriminative. In our work, we focus on gaining insights to extend the powerful class of CPT methods to purely generative UDA, where prediction on the downstream task is treated as a next word prediction task. Through this, we also present novel findings on the impact of CPT to prompt-based classifiers in the UDA framework, countering previous findings from other studies in a single-domain setting (Gu et al., 2020; Levine et al., 2021; Wettig et al., 2023).
9 Conclusion
We introduce the setting of Generative UDA, and perform an investigation on the utility of continued pre-training in this setting. We compare the approach with the popular class of domain-invariance based methods for UDA, showing that CPT is both competitive with, and more stable than invariance based approaches. Our experiments show that the benefits of CPT extend to different architectures, tuning methods and data regimes. We motivate the need for target domain exposure through CPT by showing that performance on the target domain gradually degrades with increasing masking rate. Finally, we shed light on the interplay between masking and classification performance, and how this aids UDA. Our analysis shows that in predicting masked words that are informative to the downstream task, the model implicitly learns about the downstream task, furthering the benefits of directly learning the task. Our work connects the body of UDA research with that of instruction tuning, enabling an initial step towards a wider applicability of modern language models.
Limitations
Our work presents an investigation into continued pre-training for UDA in a generative setting. Since generative UDA is an almost completely unexplored area, we establish a proof of concept by using sentence classification tasks for our analysis. We leave the extending our analysis to more complex tasks to future work.
In our study, we consider a class of PEFT methods that involve inserting learnable parameters between the layers of the model. Other classes of PEFT methods were not considered. However, we use Adapters and He et al. (2022) have shown connections between the method with Prefix Tuning  (Li & Liang, 2021) and LoRA (Hu et al., 2022).
Due to the high variance across runs in PEFT-based learning, we note that the performance can vary significantly across random seeds. We attempt to make our findings reproducible by averaging every experiment over 3 seeds. Taking environmental costs into consideration, we reduce our computational budget by running a majority of our experiments with a smaller-sized model. Learning with larger models is discussed in Section 4.
Ethics Statement
Our project aims to extend the problem of unsupervised domain adaptation to the generative setting, matching current needs with large language models. This is an effort towards improving the reliability and safety of language models, which can be fragile under distribution shift Ribeiro et al. (2020) and incur great costs over incorrect predictions Ulmer et al. (2020); Zhang et al. (2021a).
Our study does not involve any human subjects or violation of legal compliance. We do not anticipate any potentially harmful consequences to our work. As detailed in Appendix A, all of our experiments are conducted using publicly available datasets. Our code shall be released for reproducibility. Through our study and releasing our code, we hope to raise stronger research and societal awareness toward the problem of unsupervised domain adaptation in natural language processing.
Appendix A Preparation of Evaluation Benchmarks
We use two classification datasets, with 5 domains each. This results in a total of 40 pairs of source and target domains. For brevity, we include results of 24 domain pairs in the main paper, and the remaining 16 in Appendix C. For both datasets, we use the train, validation and test splits from Malik et al. (2023). More statistics about each dataset is available in Table 4. The listed datasets are intended for research purposes only. We do not make any commercial use of them.
MNLI
The Multigenre Natural Language Inference (MNLI) corpus Williams et al. (2018) contains sentence pairs across multiple genres: Travel (T), Fiction (F), Government (G), Slate (S) and Telephone (Te). The NLI task involves classifying every premise-hypothesis sentence pair as Entailment, Neutral or Contradiction.
Amazon
The Multi Domain Sentiment Analysis Dataset Blitzer et al. (2007) contains Amazon product reviews for different type of products. We use reviews from the Apparel (A), Baby (B), Books (Bo), Cameras (C) and Movies (M) domains. Each review is labelled as positive or negative.
Dataset Language License Statistics per Domain
Train Val Test
MNLI English cc-by-4.0 69600* 7735** 1945
Amazon English cc-by-4.0 1440 160 400
Table 4: Artifacts used in our study. The dataset statistics report the values used in our study.
* All domains contain approximately 69,600 examples. The exception is the Telephone domain, with 75,013 examples.
** All domains contain 7735 validation examples, except for Slate and Telephone, which contain 7731 and 8336 examples respectively.
Appendix B Details on Implementation
Models and Implementation
We use T5v1.1, T0 and GPT-2 and LLaMA-2 from the HuggingFace library [^8] https://github.com/huggingface/transformers, and use PyTorch [^9] https://pytorch.org/ to train our models.
Training
We use the default hyperparameters from Liu et al. (2022), except for batch size and training duration. We perform a grid search for these values. We train each training phase for 30,000 steps on MNLI and 15,000 steps on the Amazon dataset, with a batch size of 8. For the T5v1.1 XL and T0 models (3B parameters each), we use a batch size of 1. We train with Adam and use a learning rate of 0.003. We set the maximum sequence length to 256 tokens. We use length normalization during evaluation, as proposed by Liu et al. (2022). For each experiment, we report the mean and standard deviation across 3 runs.
[…]
Training Steps Source Accuracy Target Accuracy
5,000 93.2 (0.4) 81.9 (0.4)
10,000 93.4 (0.5) 81.6 (0.6)
15,000 93.5 (0.4) 83.3 (0.5)
[…]
Computations
Using the (IA)3 PEFT framework, training the T5v1.1 Base model (60 million parameters) for 15,000 steps takes approximately two hours on a single NVIDIA RTX A6000 GPU. The T5v1.1 XL model and T0 model (3 billion parameters) take approximately 8 hours for 15,000 steps of training. For reproducibility, each experiment is repeated thrice, with changing random seeds. In total, we run 540 experiments with the Base model and 72 experiments with the larger models. This results in a total compute time of approximately 2400 GPU hours.
Appendix C Detailed results with the Amazon and MNLI Datasets
Table 6 shows the performance of CPT on the Amazon and MNLI datasets.
On the Amazon dataset, CPT is competitive with the state of the art UDAPTER method from Malik et al. (2023) on average. We confirm this by checking for a significant difference in the performance of CPT and UDAPTER on the 20 dataset pairs. The Mann-Whitney U test and Student’s t-test both resulted in non-significant p-values of 0.5516 and 0.8316, confirming the hypothesis that there is no significant difference between CPT and UDAPTER on the Amazon dataset.
However, on the MNLI dataset, where all domains have larger gaps, both significant tests showed a significant difference between CPT and UDAPTER, with CPT being more powerful. This is exemplified through cases like Travel (T) $\to$ Government (G), where CPT yields an accuracy of 83.6% on the target domain, equalling the upper bound of the Src+Tgt baseline.
Amazon MNLI
Src+Tgt UDAPTER CPT Src+Tgt UDAPTER CPT
A $\rightarrow$ B 94.7 (0.2) 93.8 (0.3) 93.9 (0.3) T $\rightarrow$ F 77.2 (0.4) 69.7 (0.8) 74.1 (0.9)
A $\rightarrow$ Bo 94.3 (0.4) 92.5 (1.1) 90.2 (1.2) T $\rightarrow$ G 83.6 (0.7) 79.3 (0.5) 83.6 (0.3)
A $\rightarrow$ C 95.0 (0.2) 91.8 (0.5) 92.1 (0.5) T $\rightarrow$ S 72.3 (0.5) 69.6 (0.1) 70.7 (0.6)
A $\rightarrow$ M 85.8 (0.5) 81.3 (0.6) 83.3 (0.5) T $\rightarrow$ Te 77.8 (0.1) 69.4 (0.8) 76.8 (0.0)
B $\rightarrow$ A 93.4 (0.3) 93.3 (0.2) 93.4 (0.4) F $\rightarrow$ T 79.9 (0.1) 69.9 (0.2) 65.4 (0.8)
B $\rightarrow$ Bo 94.7 (0.7) 93.8 (0.3) 92.2 (0.1) F $\rightarrow$ G 82.3 (0.1) 54.3 (23.4) 78.8 (2.5)
B $\rightarrow$ C 94.7 (0.8) 93.4 (0.1) 92.1 (0.3) F $\rightarrow$ S 72.1 (0.2) 64.6 (1.8) 65.3 (1.6)
B $\rightarrow$ M 85.3 (0.2) 81.3 (0.7) 82.8 (0.2) F $\rightarrow$ Te 78.3 (0.6) 64.6 (0.7) 72.5 (0.2)
Bo $\rightarrow$ A 94.6 (0.3) 91.6 (0.5) 91.3 (0.2) G $\rightarrow$ T 79.9 (0.4) 75.9 (0.3) 75.8 (0.6)
Bo $\rightarrow$ B 94.8 (0.2) 92.9 (0.6) 90.9 (0.2) G $\rightarrow$ F 76.7 (0.1) 69.9 (0.2) 73.5 (0.2)
Bo $\rightarrow$ C 94.3 (0.2) 89.8 (0.1) 90.3 (0.4) G $\rightarrow$ S 73.1 (0.0) 69.4 (0.1) 68.0 (1.8)
Bo $\rightarrow$ M 85.5 (0.9) 84.6 (0.7) 80.1 (1.2) G $\rightarrow$ Te 78.1 (0.6) 69.9 (0.3) 73.5 (0.6)
C $\rightarrow$ A 93.4 (0.4) 92.3 (0.3) 92.5 (0.6) S $\rightarrow$ T 79.5 (0.3) 74.4 (1.7) 76.8 (0.1)
C $\rightarrow$ B 95.0 (0.6) 94.1 (0.1) 92.1 (0.2) S $\rightarrow$ F 77.7 (0.2) 73.1 (0.0) 72.4 (0.5)
C $\rightarrow$ Bo 93.9 (0.8) 91.3 (0.5) 89.0 (0.1) S $\rightarrow$ G 83.4 (0.2) 78.2 (0.5) 76.3 (0.9)
C $\rightarrow$ M 85.8 (0.1) 81.5 (0.7) 79.7 (1.2) S $\rightarrow$ Te 78.5 (0.0) 66.7 (0.2) 74.8 (0.3)
M $\rightarrow$ A 94.2 (0.7) 89.1 (1.4) 90.1 (0.5) Te $\rightarrow$ T 79.8 (0.3) 71.4 (0.0) 76.5 (0.4)
M $\rightarrow$ B 95.3 (0.5) 81.0 (16.1) 89.9 (1.2) Te $\rightarrow$ F 77.9 (0.1) 69.9 (0.5) 74.3 (0.5)
M $\rightarrow$ Bo 94.1 (0.4) 80.5 (18.6) 91.5 (0.0) Te $\rightarrow$ G 82.5 (0.1) 75.6 (1.6) 82.0 (0.6)
M $\rightarrow$ C 94.3 (0.5) 90.5 (0.0) 89.7 (0.3) Te $\rightarrow$ S 72.2 (0.0) 68.0 (0.4) 71.3 (0.5)
[…]
Figure 8: UMap visualizations of sentence embeddings from the Apparel $\to$ Movies data pair, using the T5v1.1 base model and (IA)3 PEFT method. Despite not promoting domain-invariance, CPT may be learning sentence embeddings that are separable by class labels, regardless of the domain of these sentences. The classification hyperplane for the source domain has been imagined as a solid line for illustration purposes, and its extension to the target domain is shown as a dashed line.
Method Accuracy
CPT 83.3 (0.9)
UDAPTER 81.3 (0.6)
DANN 52.3 (1.7)
CORAL 80.9 (0.4)
Task Vectors 48.0 (0.7)
Task Vectors (fine-tuning) 69.0 (0.4)
[…]
Comparison with other Model Centric Approaches
In addition to the MMD based method of Malik et al. (2023), we also compare CPT with other methods that promote domain invariance: 1) DANN (Ganin et al., 2016), which is the most widely used UDA method in NLP (Ramponi & Plank, 2020), but has been shown to be highly unstable; 2) CORAL (Sun et al., 2017), which minimizes second order statistics of the data embeddings. Additionally, with an emerging class of weight interpolation based methods, we make a comparison with task vector arithmetic Ilharco et al. (2022). The use of task vectors with PEFT methods beyond LoRA (Hu et al., 2022) has been unexplored in the literature, and we find that the method does not work with IA3. With fully fine-tuned models, the method improves in performance, but is still weaker than CPT.
CPT may learn representations that generalize across domains
[…]
Appendix D CPT across Model Architectures and Scales
We evaluate the performance of CPT over T5v1.1 XL and the instruction tuned T0 (3B) (Sanh et al., 2022) in Table 8.
Model Src+Tgt UDAPTER CPT
T5 v1.1 Base 85.8 (0.5) 78.6 (1.3) 83.3 (0.5)
T5 v1.1 XL 93.0 (0.5) 65.2 (9.5) 92.0 (1.5)
T0 3B 92.2 (0.7) 51.8 (0.8) 93.8 (0.4)
Table 8: The performance gap between CPT and UDAPTER increases with larger models, from T5v1.1 Base (60M parameters) to T5v1.1 XL (3B parameters), and further increases with instruction tuning (T0 3B).
Appendix E PEFT Frameworks
The framework proposed in Section 2 is general and can be applied to fine-tune all model parameters. Additionally, our CPT framework is compatible with the parameter-efficient fine-tuning approach. The PEFT approach is desirable because it adds only a small amount of learnable parameters $\phi$ to a pre-trained language model $\theta$, and fine-tunes only $\phi$ to perform prediction while keeping the other model parameters $\theta$ frozen. We use two instantiations in our implementations: Adapters Houlsby et al. (2019) and (IA)3 Liu et al. (2022).
(IA)3
is a state of the art PEFT learning method, and uses around a tenth of learnable parameters compared to popular methods like Adapters. (IA)3 works by element-wise multiplication (i.e. rescaling) of the model’s activations against a learned vector. In this case, the set of learnable parameters $\phi$ is a set of vectors $\{l_{\text{v}},l_{\text{k}},l_{\text{ff}}\}$ applied to each attention mechanism and feed-forward layer as,
$h$ $=\sigma\left(\frac{Q(l_{\text{k}}\mathbin{\ooalign{$\circledcirc$\cr$\bullet$}}K^{T})}{\sqrt{d_{k}}}\right)(l_{\text{v}}\mathbin{\ooalign{$\circledcirc$\cr$\bullet$}}V)$
$h$ $=(l_{\text{ff}}\mathbin{\ooalign{$\circledcirc$\cr$\bullet$}}\gamma(W_{1}x)W_{2})$
Here, $K$, $Q$ and $V$ are the key, query and value representations used in an attention block, and $W_{1}$ and $W_{2}$ are the weights in the feed-forward layer following an attention block. $l_{\text{k}}\in\mathbb{R}^{d_{k}}$, $l_{\text{v}}\in\mathbb{R}^{d_{v}}$, $l_{\text{ff}}\in\mathbb{R}^{d_{\text{ff}}}$, $\sigma$ is the softmax function while $\gamma$ is any non-linearity.
Intuitively, each vector $l$ simply learns weights measuring the importance of each feature in an activation of the pre-trained model, for the specific downstream task the model is trained on.
Adapters
are a popularly used and high performing PEFT framework, and He et al. (2022) have shown equivalence in the operations applied by Adapters, Prefix Tuning Li & Liang (2021) and LoRA Hu et al. (2022).
Adapters work by adding small learnable modules between transformer layers. Specifically, down and up projections $W_{\text{down}}\in\mathbb{R}^{d\times r}$ and $W_{\text{up}}\in\mathbb{R}^{r\times d}$ are learnt such that $\phi=\{W_{\text{up}},W_{\text{down}}\}$. A residual connection and non-linearity $\gamma$ is added at every layer,
$h=h+\gamma(hW_{\text{down}})W_{\text{up}}$
Table 9 shows CPT beats UDAPTER across different tuning methods. We also note that fine-tuning yields slightly better performance for all UDA methods.
Method Src+Tgt UDAPTER CPT
Fine-Tuning 86.4 (0.4) 82.4 (1.6) 84.4 (0.3)
(IA)3 85.8 (0.5) 81.3 (0.6) 83.3 (0.5)
Adapters 85.3 (0.5) 79.1 (0.3) 82.7 (0.5)
Table 9: Performance of CPT across different adaptation methods with the T5v1.1 base model on the Apparel $\to$ Movies domain pair. CPT remains more powerful than UDAPTER across all methods.
Appendix F CPT in a Few-Shot Setup
Table 10 accompanies Figure 2 (Section 4), showing the 256-shot performance of CPT and other baselines, across model sizes. Similarly, Table 11 accompanies Figure 4, showing the relative performance of all baselines across varying $k$.
Model Src+Tgt UDAPTER CPT
T5v1.1 Base 77.8 (0.4) 60.9 (1.6) 73.1 (1.7)
T5v1.1 XL 84.4 (0.1) 84.8 (1.5) 89.9 (1.1)
T0 3B 88.3 (0.5) 81.8 (1.3) 93.9 (0.4)
Model Src+Tgt UDAPTER Two Phase CPT
T5v1.1 Base 82.8 (0.6) 62.5 (0.7) 79.8 (1.4) 81.2 (0.7)
T5v1.1 XL 92.5 (0.4) 71.7 (7.8) 84.3 (0.9) 86.8 (2.2)
T0 3B 91.8 (0.6) 79.5 (6.7) 53.5 (0.4) 92.8 (0.2)
Table 10: Performance of CPT across different models, in a k-shot learning setup on the Apparel $\to$ Movies domain pair. We see CPT retaining strong performance on the target domain across models. Left: 32-shot. Right: 256-shot.
Number of Shots Src+Tgt UDAPTER Two Phase UDAPTER CPT
32 77.8 (0.4) 60.9 (1.6) 59.4 (2.0) 73.1 (1.7)
128 82.5 (0.5) 75.1 (0.6) 62.8 (0.6) 78.8 (1.0)
256 82.8 (0.6) 62.5 (0.7) 79.8 (1.4) 81.2 (0.7)
Table 11: Performance of CPT across different number of shots, on the Apparel $\to$ Movies domain pair, using the T5v1.1 base model. We see CPT retaining strong performance on the target domain across shots.
Appendix G Impact of Target Domain Exposure
[…]
Table 12 accompanies results from Figure 5, which show the impact of varying masking rates on CPT. Using the T5v1.1 base model, we train CPT using varying random masking rates on the Apparel $\to$ Movies domain pair, and report the mean and standard deviation over three runs. With high masking rates, the performance on the source domain is largely maintained, but the performance on the target domain rapidly deteriorates.
Masking Rate Accuracy
Source Target
5% 92.8 (0.8) 78.8 (1.8)
15% 93.5 (0.4) 83.3 (0.5)
30% 92.8 (0.6) 78.8 (1.4)
60% 92.5 (0.9) 71.0 (3.0)
90% 92.3 (0.5) 70.4 (1.5)
Table 12: Impact of Masking Rate on CPT. We train CPT using varying random masking rates on the Apparel $\to$ Movies domain pair. With high masking rates, the performance on the source domain is largely maintained, but the performance on the target domain rapidly deteriorates.
Appendix H Understanding how CPT aids UDA
Table 13 (accompanies Figure 7) shows the impact of masking sequences at inference, on classification accuracy. Words are selected for masking based on their their “informativeness”, measured by their PMI to the inference class label. The performance of the model is best with the original unmasked sequences, indicating the presence of both informative and uninformative words are essential for strong classification performance.
Table 15 accompanies Figure 7 and shows the impact of varying masking strategies on classification performance, in a few-shot setting. We also consider two different few-shot setups: one with access to the full unlabelled datasets in phase 1 pre-training, and another where even the unlabelled data is few-shot.
To isolate any effects of PEFT methods or pre-training data, we repeat the analysis from Table 15 in Table 14 with fine-tuning Flan-T5 in a full data setting, and note similar trends.
Method Accuracy
Source Target
Original 93.5 83.3
Informative Masking 88.0 78.8
Uninformative Masking 92.0 79.0
Table 13: Impact of masking at inference. We evaluate CPT on the Apparel $\to$ Movies domain pair, and select words for masking based on their “informativeness” to the classification task.
Masking Strategy Accuracy
Source Target
Random 95.8 (0.0) 86.8 (0.3)
Informative 93.9 (0.6) 85.3 (0.3)
Uninformative 95.0 (0.0) 84.8 (0.1)
Table 14: Impact of word selection for masking during training, using Flan-T5 base and no PEFT methods.
Phase 1 Data Masking Strategy Accuracy
Source Target
256 Shot Random 91.0 (0.9) 78.1 (2.4)
Informative 90.4 (0.5) 76.0 (0.7)
Uninformative 89.6 (1.2) 73.5 (1.6)
Full Data Random 90.1 (0.5) 81.2 (0.7)
Informative 91.8 (0.5) 78.0 (0.9)
Uninformative 89.3 (0.5) 72.8 (1.1)
Table 15: Impact of word selection for masking, in a 256-shot learning setup. We evaluate CPT on the Apparel $\to$ Movies domain pair, and select words for masking based on their “informativeness” to the classification task. Random masking is most powerful for the target domain, indicating that both semantic and background features are necessary for effective classification on the unlabelled domain. However, informative masking is significantly more useful than uninformative masking.
Appendix I Single Phase CPT Training
Our proposed approach in Section 2 involves two stages of training, which is more expensive than standard single phase UDA approaches. In this section, we propose a single training phase variant to CPT, and show that it performs similarly to the original method. We use the two phase pipeline in our experiments in the main paper, but note that the single and two phase pipelines are interchangeable.
We simply replace the two phase training with a joint multi-task objective as follows,
$\mathcal{L(\mathcal{D},\mathcal{D}_{\text{src}};\theta)}$ $=\frac{1}{|D|}\frac{1}{|D_{\text{src}}|}\sum_{x^{\prime}\in\mathcal{D}}\sum_{(x,y)\in\mathcal{D}_{\text{src}}}$
$(\lambda\;l(\mathbb{C}(x,y);\theta)$
$+(1-\lambda)\;l(\mathbb{M}(x^{\prime});\theta))$
where $l$ is the cross-entropy loss defined in Eq. (1), and $\mathbb{M}$ and $\mathbb{C}$ are the templates defined in Section 2. $\lambda$ is the adaptation factor which gradually changes from 0 to 1 over the course of training. This results in the model being trained almost exclusively on the MLM task early on in training, and the CLS task towards the end of training.
Table 16 compares the performance of the single phase and two phase variants of CPT. We also compare with a vanilla joint single phase objective, where $\lambda$ is fixed at 0.5 through training (called Single Phase Vanilla). The performance of the single and two phase variants are almost identical, and either can be used interchangeably. In comparison, the vanilla single phase method is significantly weaker on the target domain.
Method Accuracy
Source Target
Two Phase 93.7 (0.3) 83.3 (0.9)
Singe Phase 93.5 (0.4) 83.3 (0.5)
Singe Phase Vanilla 93.6 (0.1) 75.0 (5.7)
Table 16: Comparison of single and two-phase variants of CPT, on the Apparel $\to$ Movies domain pair. The single and two phase variants are almost identical in performance.
Appendix J Instability of Domain Invariance Methods for UDA
The Maximum Mean Discrepancy (MMD) Gretton et al. (2012) measures the difference between first order moments of variables in a Reproducing Kernel Hilbert Space Aronszajn (1950). Multiple lines of work have shown that minimizing divergence measures like MMD, when combined with auxiliary task-specific loss functions, results in training instabilities and vanishing gradients Kashyap et al. (2021); Han & Eisenstein (2019).
Figure 9: UMap visualizations of sentence embeddings from the Apparel $\to$ Movies data pair, using the T5v1.1 base model and (IA)3 PEFT method. Training with UDAPTER risks stability issues, and all embeddings from the target domain can be mapped to the closest source class cluster. This results in poor classification performance on the target domain.
We also note that as minimizing MMD does not use any label information, there is a possibility for embeddings of the target domain to be aligned with the closest source domain class cluster. For example, Figure 9 shows us a setting where both classes of the target domain (shown in green and gray) are mapped to the cluster of negative class source embeddings (shown in blue).
We compare variants of the UDAPTER method in Table 17 and show that the loss is sensitive to small changes in the loss design. Specifically we compare the UDAPTER method used in the main paper with:
  • MMD over Logits: Measures the MMD between the logits of source and target domains, instead of using intermediate model outputs.
  • Fixed Weight MMD: Instead of the multi-task loss for the MMD reduction and classification tasks, we use fixed weights for both tasks [^10] For the weighted loss, $\mathcal{L}_{\text{CLS}}+3\;\mathcal{L}_{\text{MMD}}$ was found to be the best performing..
  • Two Phase MMD: The first training phase is used to minimize MMD between source and target embeddings, while the second phase is used to train the model for classification on the source domain.
CPT remains more powerful than all variants.
Method Accuracy
Source Target
CPT 93.7 (0.3) 83.3 (0.9)
UDAPTER 94.7 (0.3) 81.3 (0.6)
UDAPTER over Logits 95.0 (0.2) 81.3 (0.7)
Fixed Weight MMD 93.4 (0.2) 78.6 (1.3)
Two Phase UDAPTER 90.1 (0.1) 68.7 (2.0)
Table 17: Comparison of variants of minimizing MMD, on the Apparel $\to$ Movies domain pair. CPT remains more powerful than all variants.